Can u solve this one plz...
1)A sum of money put out at compound interest increased by $600 at the end of the 1st year and $660 at the end of the 2nd year.Find the rate percent and the sum..

Answer by venugopalramana(3286)

Can u solve this one plz...
1)A sum of money put out at compound interest increased by??[IS THIS BY OR TO??] $600 at the end of the 1st year and $660..[THIS ALSO??] at the end of the 2nd year.Find the rate percent and the sum..
IF SUM INCREASED BY 600 IN THE I YEAR IT CAN NOT INCREASE BY ONLY 660.IT HAS TO INCREASE BY AT LEAST >600+600..SO PLEASE CHECK UP.
IF IT IS INCREASED TO THEN WE GET
LET THE SUM =P
RATE % =R
IN 1 YEAR P WILL INCREASE TO
P(1+R/100)=600.......................1
IN 2 YEARS P WILL INCREASE TO
P(1+R/100)^2=660..........................2
EQN.2/EQN.1 GIVES
1+R/100=660/600=1.1
R/100=1.1-1=0.1
R=100*0.1=10
PUTTING IN EQN.1
P*1.1=600
P=600/1.1=545.45